# Fixed-Point RNNs: From Diagonal to Dense in a Few Iterations

### v1

@@ 4.4 Dependence on $h_{t-1}$ in theory
4.5 Expressing linear attention with $\lambda_{t}\approx\mathbf{1}~{}\forall t$
Recent parameterizations of transition matrices $\lambda_{t}$ use the exponential of a negative number as opposed to the sigmoid function due to the saturation of the sigmoid (Gu et al., 2022). However, this parameterization still does not provide a mechanism to express linear attention with $\lambda_{t}\approx\mathbf{1}$ in a controlled way. To that end, De et al. (2024) propose to separate the lower-bound of $\lambda_{t}$ from its selective component $\Delta_{t}$ by setting $\lambda_{t}=\exp\left(-c\cdot\text{softplus}\left(\omega\right)\cdot\Delta_{t}\right)$. With a selective parameter $0\preceq\Delta_{t}\preceq\mathbf{1}$, an input independent component $\omega$, a temperature $c$, and the softplus function, the transition matrix $\lambda_{t}$ obeys the lower-bound $\lambda_{t}\succeq\exp\left(-c\cdot\text{softplus}\left(\omega\right)\right)$. While the selective component $\Delta_{t}$ seems to be a crucial element for certain tasks in a linear RNN (Gu & Dao, 2023), we believe it also introduces a recency bias to the recurrence. We present the following theorem as evidence for this claim.
Theorem 4.1.
Let $h_{t}=\lambda_{t}\cdot h_{t-1}+x_{t}$ define a 1-dimensional RNN parameterized by $\omega,b$ as $\lambda_{t}=\omega^{\top}x_{t}+b$. Let the input $x_{t}\sim\mathcal{N}\left(0,1\right)$, the bias term positive, and the gate be parameterized as $\lambda_{t}=\sigma\left(\omega^{\top}x_{t}+b\right)$. We define the expected memory of the RNN model $\mathcal{M}\left(\omega,b\right)$ as the expected length of sequence $\mathcal{T}$ in which we have $\lambda_{t}>0.99$ for $t\in(0,\mathcal{T}]$. Then we have: $\mathcal{M}\left(\omega,b\right)\propto\exp\left(\|\omega\|^{2}_{2}\right)^{-1}$. Proof in App. C.
Theorem 4.1 shows an exponential decrease in memory capacity as the selective component $\Delta_{t}$ becomes more prominent. This means that the weights mapping an input $x_{t}$ to $\Delta_{t}$ need to be initialized to 00 in order facilitate $\lambda_{t}\approx\mathbf{1}~{}\forall t$. In that way, the model avoids the recency bias linked to selectivity and empirically improves the performance on the copy task.
5 The Bedouin
In this section, we combine the findings from the previous sections into a dense variant of Mamba (Gu & Dao, 2023). After a short recap of Mamba, we introduce the fixed-point iteration and parametrization of our new model which we call The Bedouin. A diagram is available in Appendix, Figure 6.
5.1 Mamba: Selective SSMs
Mamba is a multi-layer network, with an embedding size of $d_{\text{model}}$. A Mamba block is a matrix state diagonal linear RNN which first expands a sequence of embeddings by a factor of $e$ to size $d_{\text{inner}}=e\times d_{\text{model}}$, and then computes an element-wise recurrence on the matrix hidden states $h_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$ as
$h_{t}=\lambda_{t}\odot h_{t-1}+\mathbf{B}_{t}\left(\Delta_{t}x_{t}\right)^{\top},$ (10)
where $\lambda_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$ is an input-dependent state transition vector, $\mathbf{B}_{t}\in\mathbb{R}^{d_{\text{state}}}$ an input transition vector, $x_{t}\in\mathbb{R}^{d_{\text{inner}}}$ the input, and $\Delta_{t}\in\mathbb{R}^{d_{\text{inner}}\times d_{\text{inner}}}$ a diagonal matrix which acts an input normalization term. The matrices are parameterized as:
$\lambda_{t}$ $=\exp\left(-\lambda_{\log}\Delta_{t}\right),$ $\lambda_{\log}$ $=\exp\left(\omega\right),$
$y_{t}^{\top}=\mathbf{C}_{t}^{\top}h_{t},\quad\mathbf{C}_{t}=\mathbf{W}_{\mathbf{C}}x_{t},$
5.2 The Bedouin iteration
Let us introduce the fixed-point iteration to the Mamba architecture. We represent the hidden state as $h_{t}^{\ell}$, where $t$ is the token index (i.e., indexing over the sequence dimension), and $\ell$ corresponds to the fixed-point iteration index (i.e., indexing over the depth dimension). The same notation is used for other variables to emphasize when they depend on the input and hidden state of the current iteration. We propose the following iteration to adapt Mamba (10) to the fixed-point mechanism for matrix state RNNs (9):
$h_{t}^{\ell}=\lambda_{t}\odot h_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\mathbf{Q}_{t}^{\ell}x_{t}\right)^{\top}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\left(\mathbf{I}-\mathbf{Q}_{t}^{\ell}\right)y_{t}^{\ell-1}\right)^{\top},$ ${y_{t}^{\ell}}^{\top}=({\mathbf{\bar{C}}_{t}^{\ell}})^{\top}h_{t}^{\ell}.$ (11)
In order to limit the Lipschitz constant according to Theorem 3.2, we use L2 normalized $\mathbf{\bar{B}}_{t}^{\ell}$ and $\mathbf{\bar{C}}_{t}^{\ell}$. Furthermore, we replace the normalization term $(\mathbf{1}-\lambda_{t})$ with $\Delta_{t}$ to stay compatible with the Mamba implementation. Expanding $y_{t}^{\ell-1}$ yields the recurrence on the matrix state
$h_{t}^{\ell}=\lambda_{t}\odot h_{t-1}^{\ell}+\text{const}+\mathbf{\bar{B}}_{t}^{\ell}(\mathbf{\bar{C}}_{t}^{\ell-1})^{\top}h_{t}^{\ell-1}(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}\Delta_{t},$
where the last term nicely illustrates the two components which mix the channels of the hidden states: the low-rank matrix $\mathbf{\bar{B}}_{t}^{\ell}(\mathbf{\bar{C}}_{t}^{\ell-1})^{\top}$ mixes over the dimension $d_{\text{state}}$, while $(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}$ mixes over the dimension $d_{\text{inner}}$. This factorization significantly simplifies the fourth-order tensor mixer formulation introduced in (8) and performs well in practice.
Finally, Bedouin (11) can be expressed as Mamba (10)
$h_{t}^{\ell}=\lambda_{t}\odot h_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\tilde{x}_{t}^{\ell}\right)^{\top},$ (12)
with an adjusted input $\tilde{x}_{t}^{\ell}=\mathbf{Q}_{t}^{\ell}\left(x_{t}-y_{t}^{\ell-1}\right)+y_{t}^{\ell-1}$. In other words, one fixed-point step consists of a channel mixing using $\mathbf{Q}_{t}$, followed by a sequence mixing using Mamba. This separation of concerns allows to speed up the parallel recurrence in time using the Mamba implementation. To find a fixed-point, the two phases are repeated until convergence, i.e. $\frac{\|y^{\ell}-y^{\ell-1}\|_{\infty}}{\|y^{\ell}\|_{\infty}}<0.1$. For a visual summary of the complete fixed-point iteration, please refer to Appendix, Figure 6.
After convergence to a fixed-point, $h_{t}^{*}$ and $y_{t}^{*}$ present the hidden state and output of the dense matrix-valued RNN. Similar to Mamba, we apply a gated linear unit to the output
$\tilde{y}_{t}^{*}=\mathbf{G}_{t}\odot\text{SiLU}\left(y_{t}^{*}\right)$
using $\mathbf{G}_{t}\in\mathbb{R}^{d_{\text{inner}}}$ and the SiLU activation function.
5.3 Parameterization
Following the analyses in Sec. 4, we propose some changes to the input-dependent parameters. Specifically, we adopt the definition of $\lambda_{t}$ from Griffin (De et al., 2024)
$\lambda_{t}$ $=\exp\left(-c\cdot\lambda_{\text{log}}\cdot\Delta_{t}\right),$ $\lambda_{\text{log}}$ $=\text{softplus}\left(\omega\right),$
$\Delta_{t}$ $=\text{diag}\left(\sigma\left(\mathbf{W}_{\Delta}^{x}x_{t}+b_{\Delta}\right)\right),$ $\mathbf{B}_{t}^{\ell}$ $=\mathbf{W}_{\mathbf{B}}^{x}x_{t}+\mathbf{W}_{\mathbf{B}}^{y}y_{t-1}^{\ell-1},$
with their proposed hyperparameter choice $c=8$, and model the dependence on the previous output with $\mathbf{C}_{t}^{\ell}=\mathbf{W}_{\mathbf{C}}^{x}x_{t}+\mathbf{W}_{\mathbf{C}}^{y}y_{t-1}^{\ell-1}$ and $\mathbf{G}_{t}=\mathbf{W}_{\mathbf{G}}^{x}x_{t}+\mathbf{W}_{\mathbf{G}}^{y}y_{t-1}^{*}$, respectively. Finally, we keep the skip connection in Mamba, but remove the short convolution due to the previous state dependency.
For the channel mixer $\mathbf{Q}_{t}^{\ell}$, we use the formulation based on the product of generalized Householder matrices and parameterize the $i=1,\ldots,r$ reflections in (6) with
$~{}\mathbf{u}_{it}^{\ell}$ $=\mathbf{W}_{\mathbf{u}_{i}}^{x}x_{t}+\mathbf{W}_{\mathbf{u}_{i}}^{y}y_{t-1}^{\ell-1},\qquad\mathbf{\bar{u}}_{it}^{\ell}=\frac{\mathbf{u}_{it}^{\ell}}{\left\|\mathbf{u}_{it}^{\ell}\right\|_{2}},$
$~{}\alpha_{it}^{\ell}$ $=\sigma\left(\left(\textbf{w}_{\alpha_{i}}^{x}\right)^{\top}x_{t}+\left(\textbf{w}_{\alpha_{i}}^{y}\right)^{\top}y_{t-1}^{\ell-1}\right).$
6 Evaluation
@@
In this section, we provide experimental results for our proposed model. For a detailed summary of the experiment setup, please refer to App. E.
@@
The task of tracking state in the alternating group on five elements ($A_{5}$) is one of the tasks introduced in (Merrill et al., 2024) to show that linear RNNs and SSMs cannot solve state-tracking problems. $A_{5}$ is the simplest subset of $S_{5}$, the word problem involving tracking the permutation of five elements. In these tasks, a model is presented with an initial state and a sequence of permutations. As the output, the model is expected to predict the state that results from applying the permutations to the initial state. Solving these task with an RNN requires either a dense transition matrix or the presence of non-linearity in the recurrence. It is therefore a good proxy to verify the state-tracking ability of Bedouin. In order to investigate the out-of-distribution generalization ability of the model, we train the model with a smaller train sequence length and evaluate for larger (more than ×$\times 3$) sequence lengths.
Figure 3: Sequence length generalization on (a) $A_{5}$ and (b) $S_{5}$. We compare a 1-layer Bedouin with a mixer $\mathbf{Q}_{t}$ of $r$ reflections to baselines with $L$ layers. The pink line denotes the train sequence length.
@@
We use the copy task (Jelassi et al., 2024) in order to assess the memory capabilities of Bedouin. In this task, the model is presented with a fixed-size sequence of elements, and expected to copy a subsequence of it after receiving a special token signaling the start of the copying process. In order to investigate the out-of-distribution generalization ability of the model, we train the models with sequence length $<50$, and assess the x2 length generalization following Jelassi et al. (2024) and Trockman et al. (2024).
Figure 4: Sequence length generalization on the copy task. We compare a 1-layer Bedouin with a mixer $\mathbf{Q}_{t}$ of $r$ reflections to baselines with $L$ layers. The gray vertical line denotes the median number of fixed-point iterations at test time. The pink line denotes the maximum train sequence length.
@@
Following Grazzi et al. (2024), we also evaluate Bedouin on the remaining unsolved task of the Chomsky Hierarchy of language problems introduced by Beck et al. (2024). Specifically, we focus on the mod arithmetic task with brackets, for which the best performance reported so far according to Grazzi et al. (2024) is an accuracy of $0.2$. Following the setup of Grazzi et al. (2024), we train on sequence lengths $3$ to $40$ and report scaled accuracies on test sequences of lengths $40$ to $256$. For Bedouin, we use a $2$-layer model with $r=4$ reflections, i.e. the best performing model in the $A_{5}$ experiment.
@@
6.1 Results
We compare Bedouin on the three tasks introduced above to Mamba (Gu & Dao, 2023), Mamba-2 (Dao & Gu, 2024), Gated DeltaNet (Yang et al., 2025) and the original LSTM (Hochreiter & Schmidhuber, 1997). In order to keep the number of layers at the same order of magnitude, we use two layers (2L) for the diagonal linear RNN baselines and one layer (1L) for Bedouin and LSTM. This keeps the number of parameters for the Bedouin comparable to the baselines (see Figure 2), and smaller than Gated DeltaNet for all experiments. For Bedouin, we report results for $r\in\{1,2,4\}$ Householder reflections of $\mathbf{Q}_{t}$ and a maximum number of $100$ fixed-point iterations. However, we also investigate the effect of limiting the number of fixed-point iterations allowed during training to $\ell_{\text{max}}\in\{2,4,8,16\}$ in Fig. 1. In order to reduce the average number of iterations, we also evaluate a randomization scheme where we sample from the Gamma distribution $\ell_{\text{max}}\sim\Gamma(4,1)$ with mean $4$ and mode $3$.
State Tracking.
In Figure 3, we compare Bedouin for $r\in\{1,2,4\}$ reflections and a maximum number of fixed-point iterations $\ell_{\text{max}}=16$ to the baselines on the $A_{5}$ and $S_{5}$ tasks with sequence length $16$. As expected, the LSTM solves $A_{5}$ and $S_{5}$, while Mamba and Mamba-2 are not able to learn it at the training sequence length.
Copying.
In Figure 4, we evaluate length generalization on the copying task . Both the Mamba and Mamba-2 models struggle with x2 generalization, which proves the effectiveness of our proposed modifications for better memory. The best-performing baseline is Gated DeltaNet, which is specifically designed to do well on the associative recall task (Yang et al., 2025) and is in fact a linear Transformer variant with about ×$\times 2$ parameters as Bedouin with mixer rank $1$. Note that the number of iterations required by Bedouin to reach the fixed point (gray vertical line) is well below the maximum sequence length of the data.
The Chomsky Hierarchy.
In Table 2, we evaluate modular arithmetic with brackets. We observe that a 2-layer Bedouin outperforms the baselines reported in (Grazzi et al., 2024) with a comparable number of parameters. In Figure 7, we plot the validation accuracy as a function of the number of fixed-point iterations. We observe that the accuracy plateaus at $20$ iterations, which is significantly less than the shortest and longest sequence in the validation set. Therefore, the number of iterations required by the Bedouin to reach its fixed point clearly does not scale with the sequence length in this task.
6.2 Required Number of Fixed-Point Iterations
A fixed-point iteration inevitably introduces sequential overhead to the computation of a model. While this might be acceptable for sequential generation at test time, reduced parallelism can be inhibiting at training time. In Figure 1, we therefore evaluate Bedouin on $A_{5}$ with a limited number of fixed-point iterations at training time $\ell_{\text{max}}$. We observe that the performance decreases once $\ell_{\text{max}}$ is lower than the training sequence length of $16$. Randomly sampling $\ell_{\text{max}}$ during training, however, allows to reduce the average sequential steps to $4$ while maintaining very high sequence length generalization at test time. We want to highlight that as explained in the previous paragraphs, we do not observe this increase in the number of fixed-point iterations in the other tasks.
7 Discussion
A fixed-point mechanism, such as the one introduced in this paper, endows a parallelizable, diagonal linear RNN with the ability to dynamically increase the sequential computation per token and describe a dense linear RNN in the limit. Our results show that such a paradigm can enable both strong state-tracking and memory capabilities with a constant number of parameters in a combined sequence and channel mixing layer (Figure 2). In fact, the fixed-point iteration gradually transforms a diagonal (i.e., channel-wise) RNN into a dense (i.e., channel-mixing) RNN, thereby allowing to trade parallel computation for expressivity (Figure 1).
For Fixed-Point RNNs to become competitive in practice, it is paramount to understand the trade-offs between parallel and sequential computation. In the worst case, if the sequential overhead is linear in the sequence length $T$, the model essentially behaves like a traditional, non-linear RNN with quadratic runtime $O(T^{2})$. This, however, is not necessarily a disadvantage if the model is capable of adapting its sequential steps to the difficulty of the task, with negligible cost for the least demanding tasks. In this paper, we focus on introducing and parameterizing the framework for Fixed-Point RNNs. Therefore, we leave the analysis and improvement of fixed-point convergence speeds beyond our preliminary results (Figure 147) to future work.
Fixed-Point RNNs present an interesting opportunity to be fused into a single GPU kernel with reduced memory I/O. This is an inherent advantage resulting from performing repeated computations on the same operands. However, in order to make the most out of this advantage, we need to consider several open problems: (1) the fixed-point needs to be independent of the level of parallelism used in fixed-point iterations, (2) the parameterization of the model needs to adhere to hardware limitations, (3) a balance between the structure and expressiveness needs to be struck when parameterizing the mixer $\mathbf{Q}_{t}$. Future progress on these problems could enable significant speed-ups in practical implementations of Fixed-Point RNNs.
Conclusion
In this paper, we presented a framework to cast a general class of dense linear RNNs as fixed-points of corresponding diagonal linear RNNs. We then show a seamless adaptation of the proposed Fixed-Point RNN framework into the linear attention family of architectures. The proposed framework provides a mechanism to trade expressivity with computation complexity while uniting the expressivity of recurrent models with the improved memory of linear attention models. Following encouraging results on toy tasks specifically designed to assess these capabilities, we hope this paper enables more expressive sequence mixing models without sacrificing memory capabilities.
Acknowledgment
Appendix
Figure 5: The difference between the fixed-point computed sequentially (i.e., computing the fixed-point for each token separately) and the fixed-point computed in parallel (i.e., computed through (11)) on the $A_{5}$ task trained on sequence length $16$ to convergence. The x-axis denotes the test sequence length, and the y-axis the normalized difference. The dashed gray line denotes the threshold for stopping the fixed-point iterations.
Figure 6: Sketch of our Beduin architecture
@@ Acknowledgment
Appendix A Proof for Theorem 3.2
We start the proof with the unrolled form of the linear RNN
$f_{\theta}(x,h)_{t}=\sum_{\tau=0}^{t}\Lambda^{t-\tau}\left(\mathbf{I}-\Lambda\right)\left(\mathbf{Q}\mathbf{B}_{\tau}x_{\tau}+\left(\mathbf{I}-\mathbf{Q}\right)h_{\tau}\right).$
Note that in order to prove the theorem, we need to show that
$\left\|\sum_{\tau=0}^{t}\Lambda^{t-\tau}\left(\mathbf{I}-\Lambda\right)\left(\mathbf{I}-\mathbf{Q}\right)(h_{\tau}-h_{\tau}^{\prime})\right\|_{2}<\left\|h-h^{\prime}\right\|_{2}.$ (13)
From the Cauchy-Schwarz inequality, we can upper-bound the LHS of (13) as
$\left\|\sum_{\tau=0}^{t}\Lambda^{t-\tau}\left(\mathbf{I}-\Lambda\right)\left(\mathbf{I}-\mathbf{Q}\right)(h_{\tau}-h_{\tau}^{\prime})\right\|_{2}\leq\left\|\sum_{\tau=0}^{t}\Lambda^{t-\tau}\right\|_{2}\cdot\left\|\mathbf{I}-\Lambda\right\|_{2}\cdot\left\|\mathbf{I}-\mathbf{Q}\right\|_{2}\cdot\left\|h_{\leq t}-h_{\leq t}^{\prime}\right\|_{2},$
where $h_{\leq t}$ corresponds to the concatenation of the hidden states $h_{\tau}$ for $\tau\leq t$. Now to prove this product is $<\left\|h-h^{\prime}\right\|_{2}$, consider the terms individually. Since $\left\|h_{\leq t}-h_{\leq t}^{\prime}\right\|_{2}\leq\left\|h-h^{\prime}\right\|_{2}$, the remaining terms need to be $<1$. Assuming $\Lambda$ is contractive, we use the Neumann series $\sum_{\tau=0}^{t}\Lambda^{t-\tau}\leq\mathbf{I}-\Lambda$ and get
$\left\|\sum_{\tau=0}^{t}\Lambda^{t-\tau}\right\|_{2}\cdot\left\|\mathbf{I}-\Lambda\right\|_{2}\leq 1.$
Finally, it remains to show that
$\left\|\mathbf{I}-\mathbf{Q}\right\|_{2}<1.$
Appendix B Proof for Theorem 3.3
Now note that from (5) we have: $\mathbf{J}_{x}^{*}\mathbf{J}_{x}^{\top}=\left(\mathbf{I}-\mathbf{J}_{h}\right)^{-1}\mathbf{J}_{x}\mathbf{J}_{x}^{\top}$. From our assumption $\mathbf{J}_{x}=\mathbf{J}_{h}:=\mathbf{J}$, we need to show that the symmetric part of the matrix $\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}\mathbf{J}^{\top}$ is positive semi-definite. Note that using the Neumann series, we can write:
Appendix C Proof for Theorem 4.1
Note that we can write $\mathcal{M}\left(\omega,b\right)$ as $\mathcal{M}\left(\omega,b\right)=\text{Prob}\left(\omega^{\top}x+b\leq\sigma^{-1}(0.99)\right)^{-1}.$ Given the distribution of $x$, we can write $\text{Prob}\left(\omega^{\top}x+b\leq\sigma^{-1}(0.99)\right)=\Phi\left(\frac{\sigma^{-1}(0.99)-b}{\|\omega\|_{2}^{2}}\right)$. Given that $\Phi(\cdot)$ can be written as the integration of a quadratic exponential function w.r.t. the denominator of its argument, this completes our proof.∎
Appendix D Low-Rank Expressiveness

### v2

@@ C.4 Dependence on $h_{t-1}$ in theory
Appendix D Evaluation
D.1 Tasks
@@
In this section, we provide experimental results for our proposed model. For a detailed summary of the experiment setup, please refer to App. D.2.
@@
The task of tracking state in the alternating group on five elements ($A_{5}$) is one of the tasks introduced in (Merrill et al., 2024) to show that linear RNNs and SSMs cannot solve state-tracking problems. $A_{5}$ is the simplest subset of $S_{5}$, the word problem involving tracking the permutation of five elements. In these tasks, a model is presented with an initial state and a sequence of permutations. As the output, the model is expected to predict the state that results from applying the permutations to the initial state. Solving these task with an RNN requires either a dense transition matrix or the presence of non-linearity in the recurrence. It is therefore a good proxy to verify the state-tracking ability of FP-Mamba. In order to investigate the out-of-distribution generalization ability of the model, we train the model with a smaller train sequence length and evaluate for larger (more than ×$\times 3$) sequence lengths.
@@
We use the copy task (Jelassi et al., 2024) in order to assess the memory capabilities of FP-Mamba. In this task, the model is presented with a fixed-size sequence of elements, and expected to copy a subsequence of it after receiving a special token signaling the start of the copying process. In order to investigate the out-of-distribution generalization ability of the model, we train the models with sequence length $<50$, and assess the x2 length generalization following Jelassi et al. (2024) and Trockman et al. (2024).
@@
Following Grazzi et al. (2024), we also evaluate FP-Mamba on the remaining unsolved task of the Chomsky Hierarchy of language problems introduced by Beck et al. (2024). Specifically, we focus on the mod arithmetic task with brackets, for which the best performance reported so far according to Grazzi et al. (2024) is an accuracy of $0.2$. Following the setup of Grazzi et al. (2024), we train on sequence lengths $3$ to $40$ and report scaled accuracies on test sequences of lengths $40$ to $256$. For FP-Mamba, we use a $2$-layer model with $r=4$ reflections, i.e. the best performing model in the $A_{5}$ experiment.
D.2 Evaluation Details
State tracking.
We train all models for $5$ epochs, with a batch size of $512$, $3$ different random seeds, learning rate set to $0.0001$, weight decay set to $0.01$, gradient clipping $1.0$, and the AdamW optimizer (Loshchilov & Hutter, 2017). For the train data, we sample $16$M datapoints from all the possible permutations for a sequence length of $16$, and split the data with a ratio of $4$ to $1$ for train and validation samples. For the test data, we sample $500$k sequences of length $50$. We use the implementation and the hyperparameters provided by Merrill et al. (2024) both for data generation and train/test. We train the model for sequence length $16$ on the train sample, and evaluate for sequence lengths $2$ through $50$ on the test sample. Consequently, each epoch of training consists of $25428$ iterations, making the total number of iterations during training to be around 1.25M. Note that the likelihood of overlap between the train and test samples is negligible since exhaustive generation of samples in $S_{5}$ and $A_{5}$ at sequence length $k$ would amount to $60^{k}$ and $30^{k}$, respectively.
Copying.
Mod arithmetic.
Language Modeling.
For the language modeling task, we use the implmentation provided by Ajroldi (2024). We use a batchsize of $16\times 4\times 4=256$, training on $4$ A100-80GB GPUs with $4$ accumulation steps, which is the batchsize used in the 2.5B setting in (Gu & Dao, 2023). The learning rate is optimized for the Mamba model ($0.004$) and train all models with this learning rate, with cosine warmup with $0.1$ steps. We use the AdamW optimizer with weight decay set to $0.1$ and $\beta_{1},\beta_{2}$ set to $0.9,0.95$.
Training Time on $A_{5}$.
In order to compare the proposed model to the baselines in terms of computation time, we train all of the baselines and our proposed model using the same hardware (A100-80GB gpus) on the $A_{5}$ task. We present the results in Fig. 6. Our Fixed-Point Mamba is trained at different maximum number of fixed-point iterations: between $2$ (green) and $16$ (blue), or sampled from the Gamma distribution $\Gamma(4,1)$ with mean $4$ (gray).
Appendix E Additional Experimental Results
E.1 Language Modeling
Figure 8: The validation perplexity of the Mamba model vs. the FPMamba model for a Householder channel mixer of $r\in\{1,2\}$ reflections. Note that all of the hyperparameters of the models are identical for fair comparison.
In order to confirm the utility of the fixed-point framework in non-state-tracking settings, we performed an experiment on language modeling. Specifically, we compare the performance of a Mamba with an FPMamba, with the same hidden size ($768$) and number of layers ($12$). The settings are selected according to the 2.5B setup introduced in Gu & Dao (2023). We use a train subsample of the FineWeb dataset (Penedo et al., 2024) with 2B tokens, and a validation subsample with 200K tokens. We use a context length of $2048$ for our experiment. For the FPMamba model, we use the Householder mixer with $1$ and $2$ reflections. We report the validation perplexity in Fig. 8.
As we can observe, the fixed-point framework does introduce a significant improvement to the performance of the model on perplexity. However, we note that this improvement cannot be only attributed to the multi-layer hypothesis of implicit models (Giannou et al., 2023), as increasing the number of Householder reflections does seem to be improving the perplexity further. Furthermore, we point out the practicality of the setup, as we can observe in Fig. 7 that in the absence of a state-tracking problem, the number of fixed-point iterations seems to be independent of the sequence length, and instead hover in the $<10$ range. Finally, fixed-point iterations are not required in the backward bass and therefore only increase training time moderately.
E.2 Long-Range State-Tracking
((a)) State Tracking on $A_{5}$ FPMamba
((b)) State Tracking on $S_{5}$ FPMamba
((c)) State Tracking on $A_{5}$ FPMamba2
((d)) State Tracking on $S_{5}$ FPMamba2
((e)) State Tracking on $A_{5}$ DeltaProduct
((f)) State Tracking on $S_{5}$ DeltaProduct
Figure 9: The state-tracking experiment for train sequence length $128$ and evaluation sequence length $[2,512]$. We omit the results of DPLR mixer due to poor performance. The figure presents (a, b) the results for FPMamba with Householder (H) and Kronecker (K) mixer, (c, d) the results for FPMamba2 with Householder (H) and Kronecker (K) mixer, and (e, f) the DeltaProduct method (Siems et al., 2025) for Householder mixers.
In this section, we investigate the ability of our proposed method in doing state tracking on longer sequences. Specifically, we will use the $A_{5}$ and $S_{5}$ datasets and train on sequence length 128, while evaluating for sequence lengths in the range $[2,512]$. We also implement the proposed Fixed-Point framework on Mamba2 (Dao & Gu, 2024), and we compared our method to DeltaProduct (Siems et al., 2025). In Fig. 9, we plot the test accuracy for these models.
Comparing our results to DeltaProduct, we can see that the non-linearity introduced by the Fixed-Point dynamics allow for a slight improvement in the performance of the Householder products as the mixer components. Furthermore, we observe that the best performing mixer variant is still the Kroneckers model, which can successfully learn the state-tracking problem in all runs. Moreover, the FPMamba2 model demonsterates a better length generalization ability compared to FPMamba1, which we attribute to the improved underlying architecture used in the model. As shown in (Dao & Gu, 2024), Mamba2 has better recall capabilities, which can help with length generalization.
E.3 Sequential vs Parallel Fixed-Point Iteration
An important detail about the fixed-point framework proposed in this paper is that it is not convex. Therefore, the fixed-point is not necessarily unique, which can be problematic in autoregressive applications because there will be no guarantee that the parallel fixed-point during training will be the same as the sequential fixed-point used during inference (Schöne et al., 2025). In order to investigate this issue, we trained an FPMamba model on the $A_{5}$ task and compared the fixed-point computed sequentially and in parallel. We report the results in Fig. 10. We observe that the fixed-points are extremely similar, providing the possiblity of computing the fixed-point sequentially during inference.
Figure 10: The difference between the fixed-point computed sequentially (i.e., computing the fixed-point for each token separately) and the fixed-point computed in parallel (i.e., computed through Eq. 12) on the $A_{5}$ task trained on sequence length $16$ to convergence. The x-axis denotes the test sequence length, and the y-axis the normalized difference. The dashed gray line denotes the threshold for stopping the fixed-point iterations.
E.4 Modular Arithmetic Task Results
In Tab. 2, we evaluate modular arithmetic with brackets. We observe that a 2-layer FP-Mamba outperforms the baselines reported in (Grazzi et al., 2024) with a comparable number of parameters. In Fig. 11, we plot the validation accuracy as a function of the number of fixed-point iterations. We observe that the accuracy plateaus at $20$ iterations, which is significantly less than the shortest and longest sequence in the validation set. Therefore, the number of iterations required by FP-Mamba to reach its fixed point clearly does not scale with the sequence length in this task.
@@ E.4 Modular Arithmetic Task Results
Appendix F Low-Rank Expressiveness